OpenStudy (anonymous):

The volume of a gas is 605 liters at 27.0°C. The new temperature is -3.0°C. What is the new volume? (Remember to convert temperatures to the Kelvin scale.)

OpenStudy (imstuck):

You first have to convert the C to K or else your answer will be a - volume and that's just not possible. The conversion is the temp C + 273.15 = temp K. So convert both your temps to K. 27C = 300.15K; -3C = 270.15K. Now set up a ratio:\[\frac{ Volume _{1} }{ Temp _{1} }=\frac{Volume _{2} }{ Temp _{2} }\]or...

OpenStudy (imstuck):

\[\frac{ 605 }{ 300.15 }=\frac{ x }{ 270.15 }\]Solve this for x to find your new volume. What do you get?

OpenStudy (anonymous):

672 L?